OpenStudy (anonymous):

Calc, help! How do I determine at what points y=x+2/cos x is continuous?

OpenStudy (anonymous):

where the function is defined

OpenStudy (anonymous):

what's the def of continuous?

OpenStudy (anonymous):

not at asymptote

OpenStudy (anonymous):

its not defined in text so yeah im assuming

OpenStudy (anonymous):

Definition in terms of limits of functions The function f is continuous at some point c of its domain if the limit of f(x) as x approaches c through the domain of f exists and is equal to f(c).[2] In mathematical notation, this is written as \[ \lim_{x \to c}{f(x)} = f(c).\] In detail this means three conditions: first, f has to be defined at c. Second, the limit on the left hand side of that equation has to exist. Third, the value of this limit must equal f(c).

OpenStudy (anonymous):

as x approaches what value is the function undefined

OpenStudy (anonymous):

take the derivative

OpenStudy (anonymous):

can you walk me through that step by step if possible? I missed the lecture and am somewhat lost

OpenStudy (anonymous):

you need to use the quotient rule and the derivative of cos is -sin

OpenStudy (anonymous):

f(x)/g(x)=l/m?

OpenStudy (anonymous):

the function is not defined when \[x=\frac{ \pi }{ 2 }+k\pi\text{, where }k \in \mathbb{Z}\]

OpenStudy (anonymous):

because cos x will be 0 at those values of x and the function will not be deifned there. thus, the function will be discontinuous at those points

OpenStudy (anonymous):

so what you have above pi/2+ k(pi), how did you determine points off that?

OpenStudy (anonymous):

y = x is continuous for all real x. y = 2/x is continuous for all real x where x is not 0. -1 <= cos x <= 1 for all real x. thus, so long as cos x not = 0, your function will be continuous.

OpenStudy (anonymous):

ohhhhh. Ok I got it now. Thank you!